# Bootstrapping Vision-language Models for Self-supervised Remote Physiological Measurement

### v1

@@ 6.2 Vision and text encoders
6.3 Losses
Frequency contrastive loss. Frequency contrastive loss $L_{fc}$ is the most important loss in existing self-supervised remote physiological measurement methods [20, 21, 22, 23]. It enforces the signal frequency similarities among positive samples and dissimilarities between positive and negative samples in the feature space. In Table VII we use VL-phys w/o $L_{fc}$ to denote a variant without using $L_{fc}$ in our framework. We can see that the MAE and RMSE are substantially increased while $\rho$ is substantially decreased from the original framework, e.g. , +2.42, +2.83 and -0.09 on MMVS, which proves the benefits of adding $L_{fc}$.
Frequency ranking loss. We verify the effectiveness of the proposed frequency ranking loss $L_{fr}$ (Sec. 3.6). First, as shown in Table VII, omitting $L_{fr}$ leads to deteriorated results on both datasets. Next, $L_{fr}$ optimizes the relative frequency ranks of predicted positive and negative signals rather than their absolute frequency differences to avoid overfitting. If we instead constrains the signal frequency ratios between these signals to be consistent with the absolute ground truth ratios $R$, denoted by $L_{fr}\rightarrow L_{afr}$ in Table VII, the performance gets worse, e.g. , +0.05 and +0.09 in MAE on two datasets, respectively.
@@ 6.4 Text prompt design
We analyze the influence of text prompt design in Table VIII. We denote VL-phys-T1, VL-phys-T2, VL-phys-T3 as three variants, each uses an alternative prompt. We observe that these varints show comparable performance to the original VL-phys, indicating that our approach remains effective as long as the text prompt is grammatically correct and can accurately depict the relative frequency ratios of skin color temporal variations in C-STMaps. This robustness highlights our method’s ability to adapt to various text prompts while consistently capturing frequency-related attributes for rPPG estimation.
Additionally, we introduce VL-phys-Video as another variant without generating spatio-temporal maps (STMaps) from positive and negative video samples. In this variant, we concatenate the positive and negative video along the temporal dimension to obtain a cross-sample video. We then define its corresponding text prompt using template shown in Table VIII. We can see the MAE respectively increases to 0.57 and 3.01 on two datasets by this variant. This can be owed to the potential noises and distractions when directly using the original videos. In contrast, STMaps retain only pixels from detected facial ROIs while masking other irrelevant pixels in the videos. They can effectively highlight the temporal variation of skin color.
TABLE IX: Zero-shot HR estimation performance of pre-trained VLMs.
@@
TABLE X: Ablation study on visual embeddings for rPPG estimation.
@@
6.5 Zero-shot rPPG estimation using pretrained VLMs
We investigate the zero-shot rPPG estimation performance using off-the-shelf pre-trained VLMs (CLIP, BLIP, ALBEF) corresponding to the discussion in Sec. 1. Initially, we define a frequency bin [0.5, 1.0, 1.5, 2.0, 2.5] and create text prompts to describe the frequency of skin color temporal variation. The text template is defined as ”the frequency of skin color temporal variation is [ratio] hertz in the video.” We respectively replace the [ratio] token with each frequency value in the bin and then calculate the similarity between the embeddings of the text prompt and the given video. After calculating the similarity score for each frequency value, we linearly combine them to obtain the final prediction value. We then multiply the results by 60111The frequency value should be multiplied by 60 to convert it from hertz to HR in beats per minute (bpm). to calculate the HR for performance comparison with our VL-phys. As shown in Table IX, we can see a substantial performance gap between using VL-phys and off-the-shelf pre-trained VLMs, highlighting their difficulties in capturing periodic visual patterns. These findings also reveal that VLMs need to be re-trained or fine-tuned based on specific vision-text pairs, otherwise they would struggle to understand the attributes related to frequency.
6.6 Visual embeddings for rPPG estimation
@@
6.7 Number of negative samples
We vary the number of negative video samples in this session. In Fig. 7, we vary this number $k$ from 1 to 6 and observe that the performance appears to be stable when $k\geq 3$. Considering that a big $k$ would increase the computation cost, we select $k=3$ as our default setting.
@@ 7 Conclusion
This paper presents a novel self-supervised learning framework that bootstraps vision-language models (VLMs) for remote physiological measurement. It develops both generative and contrastive learning mechanisms to enhance the VLM’s ability to capture the frequency of skin color temporal variation for rPPG estimation. For a given facial video, our VL-phys begins by creating positive and negative spatio-temporal maps (STMaps) with varying rPPG signal frequencies. Then we horizontally concatenate them to create contrastive spatio-temporal maps (C-STMaps). We also construct text prompts to describe the relative ratios of signal frequencies across different parts of C-STMaps; this forms frequency-oriented vision-text pairs. Next, we fine-tune the pre-trained vision and text encoders of VLM using these vision-text pairs via multimodal generative and contrastive tasks. These include the text-guided visual map reconstruction task aimed at recovering masked image patches of C-STMaps with the guidance of text prompts, and the vision-text contrastive learning task to align vision and text modalities. Moreover, we also propose the unimodal frequency contrastive and ranking task to optimize the estimated rPPG signals across multiple video samples. Extensive experiments on four datasets demonstrate that our method not only outperforms existing self-supervised methods but also rivals state of the art supervised methods in HR, HRV, and RF estimations.

### v2

@@ 6.4 Text prompt design
We analyze the influence of text prompt design in Table 9. We denote VL-phys-T1, VL-phys-T2, VL-phys-T3 as three variants, each uses an alternative prompt. We observe that these varints show comparable performance to the original VL-phys, indicating that our approach remains effective as long as the text prompt is grammatically correct and can accurately depict the relative frequency ratios of skin color temporal variations in C-STMaps. This robustness highlights our method’s ability to adapt to various text prompts while consistently capturing frequency-related attributes for rPPG estimation.
Additionally, we introduce VL-phys-Video as another variant without generating spatio-temporal maps (STMaps) from positive and negative video samples. In this variant, we concatenate the positive and negative video along the temporal dimension to obtain a cross-sample video. We then define its corresponding text prompt using template shown in Table 9. We can see the MAE respectively increases to 0.57 and 3.01 on two datasets by this variant. This can be owed to the potential noises and distractions when directly using the original videos. In contrast, STMaps retain only pixels from detected facial ROIs while masking other irrelevant pixels in the videos. They can effectively highlight the temporal variation of skin color.
Table 10: Zero-shot HR estimation performance of pre-trained VLMs.
@@
6.5 Zero-shot rPPG estimation using pretrained VLMs
We investigate the zero-shot rPPG estimation performance using off-the-shelf pre-trained VLMs (CLIP, BLIP, ALBEF) corresponding to the discussion in Sec. 1. Initially, we define a frequency bin [0.5, 1.0, 1.5, 2.0, 2.5] and create text prompts to describe the frequency of skin color temporal variation. The text template is defined as ”the frequency of skin color temporal variation is [ratio] hertz in the video.” We respectively replace the [ratio] token with each frequency value in the bin and then calculate the similarity between the embeddings of the text prompt and the given video. After calculating the similarity score for each frequency value, we linearly combine them to obtain the final prediction value. We then multiply the results by 60111The frequency value should be multiplied by 60 to convert it from hertz to HR in beats per minute (bpm). to calculate the HR for performance comparison with our VL-phys. As shown in Table 10, we can see a substantial performance gap between using VL-phys and off-the-shelf pre-trained VLMs, highlighting their difficulties in capturing periodic visual patterns. These findings also reveal that VLMs need to be re-trained or fine-tuned based on specific vision-text pairs, otherwise they would struggle to understand the attributes related to frequency.
6.6 Visual embeddings for rPPG estimation
[…]
Table 11: Ablation study on visual embeddings for rPPG estimation.
@@ 6.6 Visual embeddings for rPPG estimation
Table 12: Ablation study on the negative sample generation method.
Method UBFC-rPPG MMVS
@@
6.7 Generation of negative samples
[…]
Moreover, we evaluate the effectiveness of our predefined frequency ratio bin used in the LFA module for negative samples generation. According to [66], the frequency of skin color temporal variation in real facial videos typically lies within the range of [0.75, 4] Hz. Based on this observation, we predefine a frequency ratio bin with a narrow value range to avoid significant frequency changes between the generated positive and negative samples. This design enables the subsequent vision encoder to focus on distinguishing the samples based on their subtle frequency differences, which is essential for accurately extracting underlying rPPG signals. If the frequency ratio $r$ is instead randomly sampled from a broader range (0, 10), as shown in the variant VL-phys (br) in Table 12, the MAE and RMSE increase significantly, e.g. +1.32 and +1.05 on the MMVS dataset. These results underscore the effectiveness and superiority of our predefined ratio bin.
We further vary the number of negative video samples in this session. In Fig. 8, we vary this number $k$ from 1 to 6 and observe that the performance appears to be stable when $k\geq 3$. Considering that a big $k$ would increase the computation cost, we select $k=3$ as our default setting.
@@ 6.7 Generation of negative samples
Table 13: Ablation study on the spatio-temporal map generation methods.
Method UBFC-rPPG MMVS
[…]
VL-phys 0.28 0.60 0.99 2.33 3.76 0.95
VL-phys(POS) 0.30 0.63 0.99 2.32 3.76 0.95
VL-phys(CHROM) 0.31 0.63 0.99 2.37 3.78 0.95
6.8 Generation of STMaps
In our framework, we use the spatio-temporal map generation method proposed in [36] and [32] to generate the STMaps for positive and negative samples. However, different kinds of STMaps can be generated by using alternative methods discussed in [28]. Table 13 presents two variants, VL-phys (POS) and VL-phys (CHROM), where our generated STMaps are replaced by POS-STMaps and CHROM-STMaps, respectively. POS-STMap consists of U and V channels derived from the original facial videos, along with a POS channel generated from the POS algorithm [57]. Similarly, CHROM-STMap also incorporates the U and V channels, but the POS channel is replaced by the CHROM channel generated from the CHROM algorithm [7]. We can see these variants have similar performance with the original VL-phys. This indicates that the proposed framework is robust and generalizable across different spatio-temporal map generation methods.
@@ 7 Conclusion
This paper presents a novel frequency-centric self-supervised learning framework that bootstraps vision-language models (VLMs) for remote physiological measurement. It develops both generative and contrastive learning mechanisms to enhance the VLM’s ability to capture the frequency of skin color temporal variation for rPPG estimation. For a given facial video, our VL-phys begins by creating positive and negative spatio-temporal maps (STMaps) with varying rPPG signal frequencies. Then we horizontally concatenate them to create contrastive spatio-temporal maps (C-STMaps). We also construct text prompts to describe the relative ratios of signal frequencies across different parts of C-STMaps; this forms frequency-oriented vision-text pairs. Next, we fine-tune the pre-trained vision and text encoders of VLM using these vision-text pairs via frequency-related multimodal generative and contrastive tasks. These include the text-guided visual reconstruction task aimed at recovering masked image patches of C-STMaps with the guidance of text prompts, and the vision-text contrastive learning task to align vision and text modalities. Moreover, we also propose the unimodal frequency contrastive and ranking task to optimize the estimated rPPG signals across multiple video samples. Extensive experiments on four datasets demonstrate that our method not only outperforms existing self-supervised methods but also rivals state of the art supervised methods in HR, HRV, and RF estimations.
Declarations
  • Competing interests
    The authors declare that they have no competing interests.
  • Data availability
    UBFC-rPPG dataset is available at https://sites.google.com/view/ybenezeth/ubfcrppg.
    PURE dataset is available at https://www.tu-ilmenau.de/universitaet/fakultaeten/fakultaet-informatik-und-automatisierung/profil/institute-und-fachgebiete/institut-fuer-technische-informatik-und-ingenieurinformatik/fachgebiet-neuroinformatik-und-kognitive-robotik/data-sets-code/pulse-rate-detection-dataset-pure.
    VIPL-HR dataset is available at https://vipl.ict.ac.cn/resources/databases/201811/t20181129_32716.html.
    MMVS is available at https://github.com/yuezijie/HR-estimation/tree/master.